### v2

# A Novel Hybrid Approach to Contraceptive Demand Forecasting: Integrating Point Predictions with Probabilistic Distributions

Abstract
Accurate demand forecasting is vital for ensuring reliable access to contraceptive products, supporting key processes like procurement, inventory, and distribution. However, forecasting contraceptive demand in developing countries presents challenges, including incomplete data, poor data quality, and the need to account for multiple geographical and product factors. Current methods often rely on simple forecasting techniques, which fail to capture demand uncertainties arising from these factors, warranting expert involvement. Our study aims to improve contraceptive demand forecasting by combining probabilistic forecasting methods with expert knowledge. We developed a hybrid model that combines point forecasts from domain-specific model with probabilistic distributions from statistical and machine learning approaches, enabling human input to fine-tune and enhance the system-generated forecasts. This approach helps address the uncertainties in demand and is particularly useful in resource-limited settings. We evaluate different forecasting methods, including time series, Bayesian, machine learning, and foundational time series methods alongside our new hybrid approach. By comparing these methods, we provide insights into their strengths, weaknesses, and computational requirements. Our research fills a gap in forecasting contraceptive demand and offers a practical framework that combines algorithmic and human expertise. Our proposed model can also be generalized to other humanitarian contexts with similar data patterns.
keywords:
family planning supply chain , hybrid forecasting , forecast distributions , contraceptive demand , forecast combination
journal: International Journal of Production Research
[…]
longtable
[…]
1 Introduction
A fundamental aspect of ensuring reliable access to contraceptive products lies in accurate demand forecasting, as demand forecasting forms the foundation of efficient and reliable procurement, sourcing, storage, allocation, and distribution processes for contraceptive products [3]. However, the task of producing accurate and reliable demand forecasts for contraceptives in developing countries presents numerous challenges [18]. These challenges include the unavailability of comprehensive data [43], poor data quality [19], and the necessity to forecast across multiple geographical and product hierarchies [62].
Despite the complexity inherent in demand forecasting, many forecasts in practice are often generated using simple methods such as the moving average of historical consumption data or demographic forecasting techniques [73]. However, these methods rarely consider the complexities introduced by users switching from one contraceptive method to another, driven by factors such as the introduction of new products, health concerns, or issues with accessibility and availability [2]. This inadequacy contributes to inefficiencies in the family planning supply chain (FPSC) affecting the availability of contraceptive products [49].
Evidence from the PMA2020 survey [^1] Performance Monitoring and Accountability 2020 survey. further underscores this issue, revealing that many health sites and contraceptive outlets in developing countries often face stockouts of contraceptive methods [1]. Such stockouts limit access to contraceptive products for users when needed, either by restricting the availability of preferred methods or by turning away users due to product unavailability [50]. Consequently, these challenges contribute to an increase in unmet demand222unmet demand is defined as the percentage of women of reproductive age who currently have a need for family planning but are either not using any contraceptive methods or whose partners are not using them [31]. for contraceptive products [8].
The unmet demand for contraceptives is a significant concern, as it leads to an estimated 121 million unintended pregnancies each year, roughly 331,000 per day [72]. This situation incurs substantial costs, both for women and children and for society at large [62]. Over 60% of unintended pregnancies end in abortion, whether safe or unsafe, legal or illegal, posing significant risks to women’s lives [10]. Unfortunately, over 45% of these abortions are unsafe and result in maternal deaths [61]. This situation is particularly worse in developing countries, where approximately 7 million women are hospitalized each year due to unsafe abortions [64]. Moreover, this also creates a public health crisis, costing an estimated 2.8 billion USD per year for abortion and post-abortion care in low- and middle-income countries [65]. Recognizing its importance, this issue has been prioritized as essential for achieving the 2023 Sustainable Development Goals (SDGs).
Despite efforts by governments, foundations, and donors to increase the uptake of contraceptive products through policy and program interventions [49], developing countries continue to experience high unmet demand, particularly due to persistent stockouts at local health sites and contraceptive outlets [63]. A key reason for this ongoing issue is that these efforts and assessments are largely focused on the national or global level, which can mask the ground reality due to local disparities [50].
Recognizing the need for an improved forecasting process at the local health site level, the United States Agency for International Development (USAID) launched the “Intelligent Forecasting Challenge: Model Future Contraceptive Use” [74]. This competition aimed to source new solutions, test novel ideas, and scale effective approaches for contraceptive demand forecasting using not only time series methods but also data driven methods like machine learning (ML). However, the competition missed a critical element of the contraceptive demand forecasting process: quantifying and communicating the uncertainty, as it focused exclusively on point forecasts. On the other hand, the FPSC in developing countries is often associated with numerous uncertainties, including complex patterns of demand, variable lead times, and dependence on donor support [48]. These factors further exacerbate demand uncertainty, necessitating the use of probabilistic estimations to quantify the uncertainty of future demand.
Discussions with USAID officials highlighted that decision-makers are particularly interested in the upper bounds of prediction intervals, as they are keen to mitigate the risk of stockouts, a critical issue in contraceptive supply chains. As LaCroix et al. [43] has noted, despite the acknowledged need for probabilistic approaches, point predictions remain the default due to the lack of standardized methodologies for incorporating uncertainty into contraceptive demand forecasts. Probabilistic estimations, by quantifying the uncertainty inherent in these predictions, thus represent a valuable tool for managing stock levels and reducing the risk of supply shortages.
To our knowledge, no previous work has focused on probabilistic forecasting in contraceptive demand estimation within the FPSC at the local healthcare site level. Thus, this paper first addresses this gap by investigating a probabilistic forecasting approach for estimating demand for contraceptive products using data from January 2016 to December 2019, extracted from the Logistics Management Information System (LMIS) of Cote d’Ivoire. This is the same dataset that was used in the competition. Moreover, since the publication of the Contraceptive Forecasting Handbook, which focuses on simple forecasting methods (see USAID [73] for more information), there has been no literature evaluating the applicability and usability of different forecasting methods in contraceptive demand forecasting. Therefore, in this study, we employ a range of forecasting methods, including time series, Bayesian, ML and foundational time series methods [^3] Foundational Time Series Methods: A class of pre-trained machine learning models designed specifically for time series forecasting, leveraging large-scale datasets to generalize across different forecasting tasks. These models require no or minimal tuning and can handle complex temporal patterns efficiently., to produce point forecasts along with probabilistic forecasts for all products across all healthcare sites.
Additionally, demand planners widely apply judgmental adjustments to incorporate external factors based on their expertise in the FPSC setting [3]. Our discussions with USAID professionals revealed similar insights; they explained that site-level demand planners often adjust system-generated forecasts or use judgmental forecasts to eliminate data inaccuracies. These inaccuracies may arise because the data used to prepare these forecasts may not reflect true demand due to stockouts or incomplete data collection, or because planners have additional information, such as product discontinuation [19]. Hence, the human factor is valuable in this forecasting setup [43].
Given that system-generated forecasts and human forecasts offer distinct benefits, it is vital to design a hybrid intelligence system that combines them. In this context, where site-level demand planners produce point forecasts, we are particularly interested in how to combine point forecasts with probabilistic forecasts to produce combined probabilistic forecasts. However, the literature often treats point forecast combination methods and probabilistic forecast combination methods separately [77]. To address this gap, we propose a Constrained Quantile Regression Averaging (CQRA) method [^4] CQRA: A statistical method that combines multiple quantile regression models while enforcing constraints to ensure coherence and interpretability of predictions. It enhances robustness by optimizing weights assigned to different models under quantile-specific constraints, improving accuracy in heterogeneous data settings. to combine point forecasts made by experts with probabilistic forecasts generated by a system-based forecasting method. We compare the forecast performance using the Mean Absolute Scaled Error (MASE), a scale-independent metric that provides robustness and stability for point forecasts, and Continuous Ranked Probability Scores (CRPS), a widely used metric in probabilistic forecasting that assesses the sharpness and calibration of the forecast distribution for probabilistic forecasts in a cross-validation setup. Finally, we compare our method results against the submissions from the competition.
Thus, our contributions are as follows:
  1. 1.
    We produce point forecasts along with forecast distributions for contraceptive products at the healthcare site level, quantifying uncertainties in future contraceptive demand.
  2. 2.
    We develop a novel method to combine point forecasts with probabilistic forecasts, allowing human experts to incorporate their expertise into the forecast and thereby providing a hybrid intelligence system.
  3. 3.
    We provide a detailed comparison of the performance of time series, Bayesian, ML and foundational time series methods, and our proposed hybrid method. Additionally, we provide a comparison of the computational requirements for each method, offering a holistic view of the differences between these forecasting methods.
  4. 4.
    We have made the code and data for our proposed method, along with all other methods used in this study, publicly accessible to ensure the reproducibility. Furthermore, our study adheres to the replication principle [12], allowing for the method’s generalization across various sectors with similar data patterns.
The remainder of the paper is structured as follows: Section 2 provides a brief overview of the literature and discusses its limitations to position our work. In Section 3, we discuss the data and the experimental setup. Section 4 presents the results of our analysis. In Section 5, we summarize our findings, discuss the limitations, and present ideas for future research directions.
2 Research background
Over the past few decades, reducing the unmet need for contraceptives has been a central focus in the field of FPSC [49]. This issue has been recognized as a critical agenda item in achieving the 2030 Sustainable Development Goals, particularly in expanding access to contraception to ensure universal access to family planning services [36]. Consequently, accurate and reliable demand forecasting plays a crucial role in the FPSC, as it supports informed decision making processes to ensure access to safe and effective contraceptives, thereby empowering individuals and communities to make informed reproductive health choices [1].
The majority of literature on forecasting in the FPSC has centered on estimating family planning indicators at national or global levels to guide strategic decisions. For instance, Ahmed et al. [1] employed linear and quadratic logistic regression methods to estimate the modern contraceptive prevalence rate (mCPR) [^5] The estimation of the percentage of women using a modern contraceptive product [1]. in five sub-Saharan African countries, using data from the PMA2020 survey. Similarly, New et al. [50] examined trends in three family planning indicators; 1) mCPR, 2) unmet demand for modern contraceptives, and 3) demand satisfied by modern contraceptives. Their study covered the period from 1990 to 2030 for 29 states and union territories in India. To conduct their analysis, they employed a Bayesian hierarchical method, integrating statistical time-series techniques with demographic factors drawn from the Demographic Health Survey (DHS) [76], Annual Health Survey, and District-Level Household Survey. Haakenstad et al. [31] used a spatio-temporal Gaussian process regression method to estimate mCPR, method mix, and demand satisfied for the global contraceptive prevalence rate between 1970 and 2019.
These national and global-level studies provide valuable strategic insights into contraceptive coverage and trends. However, they were not designed to address operational realities at local levels, such as variations in demand and supply chain issues at individual service delivery points. Their focus remains on broader, aggregate-level decisions that inform national or global policies and strategies [50]. Thus, while these studies may not capture the granular, site-specific challenges at the operational level, their contributions are still highly valuable for higher-level planning and resource allocation.
On the other hand, a fewer have addressed national-level forecasts for specific contraceptive products. For instance, Akhlaghi et al. [2] estimated national demand for condoms using demographic-based forecasting and a consumption-based moving average method, incorporating expert judgment to refine predictions. Karanja et al. [37] applied consumption-based forecasting using Auto Regressive Integrated Moving Average (ARIMA) and exponential smoothing methods to estimate demand for contraceptive pills, injectables, implants, and intrauterine contraceptive devices (IUDs), using data from Kenya’s District Health Information System (DHIS). Moreover, Khan et al. [39] used a demographic-based forecasting approach integrating expert assumptions with DHS data to estimate demand for Sayana Press, a new injectable contraceptive, across 12 countries. These studies also reflect a focus on national or strategic decision-making.
However, at the operational level, where service delivery occurs, forecasting must account for uncertainties and variations specific to each location. In contrast to aggregate studies that often incorporate probabilistic forecasting, which includes prediction intervals to manage uncertainty [50, 1, 31], operational studies like Akhlaghi et al. [2] and Karanja et al. [37] primarily used point forecasts. These point forecasts provide a single estimate without explicitly addressing the uncertainty surrounding the forecasted values. In dynamic settings such as healthcare service delivery, this can limit their applicability. Operational-level contexts often experience demand variations due to factors like stockouts, local preferences, and seasonal changes [49].
Probabilistic forecasts, which include a range of possible outcomes (such as prediction intervals), are crucial in such contexts because they acknowledge the uncertainty and help supply chain managers make more informed decisions [60]. For example, Khan et al. [39] utilized scenario analysis to acknowledge forecast uncertainty for the injectable contraceptive Sayana Press in 12 countries. However, in the works of Akhlaghi et al. [2] and Karanja et al. [37], only point predictions were generated, with no explicit consideration of uncertainty in the forecasts. This omission potentially limits the applicability of their findings in dynamic operational contexts, as point forecasts, which provide a single value estimate, are simpler but less adaptable to fluctuating demand conditions.
In practice, however, many still rely on simple methods to produce point forecasts, despite the limitations these approaches present in capturing the demand variations for each contraceptive product. [3]. These methods often fail to address the complexity of demand and attempt to answer multiple questions using point forecasts [43]. The most commonly employed methods include: 1) extrapolating historical consumption using basic time series methods, linear trends, averages, or simple regression methods; 2) estimating consumption based on service statistics, such as program plans; and 3) utilizing population demographics to project demand [73]. In practice, however, many supply chain managers still rely on point forecasts due to their simplicity, despite the limitations of these approaches in accounting for demand fluctuations at the local level [3]. Common methods include basic time series methods, linear trends, and demographic-based projections [73]. While these methods are easy to implement, they often overlook critical uncertainties and the complexities of real-world contraceptive demand, such as external shocks (e.g., global crises like COVID-19), demand shifts due to market cannibalization, or supply chain disruptions [43].
As a result, these simplistic forecasting approaches can lead to inefficiencies in ordering and distribution, causing stockouts or overstocking at healthcare sites, ultimately impairing access to essential contraceptive services [49]. This underscores the need for more advanced forecasting methods that incorporate both uncertainty and local variations in demand to optimize supply chain performance and support better family planning outcomes [8].
2.1 USAID Intelligent Forecasting Competition
The USAID Intelligent Forecasting Competition attempted to address the need for a more reliable contraceptive demand forecasting method by inviting participants to develop intelligent forecasting methods. Specifically, participants were tasked with forecasting contraceptive consumption at the service delivery level of Côte d’Ivoire’s public sector health system over a three-month forecast horizon, using data provided from the Cote d’Ivoire public health system to forecast the consumption of contraceptives over three months. This competition attracted nearly 80 submissions from 40 participants, reflecting a diverse range of forecasting approaches [74].
The winning entry in the USAID Intelligent Forecasting Competition, developed by Inventec Corporation, used distinct models for each forecasting horizon. They employed both LightGBM and Long Short-Term Memory (LSTM) methods, incorporating categorical features, historical data, and future population projections. The final forecast was derived from an ensemble of these methods, with weights assigned differently for each forecasting horizon.
The second-place model employed an ensemble approach, averaging the predictions from six separate LightGBM models, each designed for a specific forecasting horizon. This strategy reflects a robust method for improving forecast accuracy through aggregation. The third-best model also used LightGBM but incorporated hyperparameter tuning and additional trend indicators, such as linear and polynomial functions, alongside various time-series features.
Another notable submission employed hierarchical forecasting with a bottom-up approach, using ARIMA (AutoRegressive Integrated Moving Average) as the base forecasting method. This method demonstrates the utility of hierarchical techniques in complex forecasting scenarios. Additionally, a model that combined neural networks with a naive forecasting approach was also among the top contenders. For longer time series, this model used an ensemble of predictions from different neural network architectures, including Convolutional Neural Networks (CNN), Gated Recurrent Units (GRU), and LSTMs, with the final prediction being the median of these forecasts. For shorter time series, it used a naive Bayes method. Table LABEL:tbl-literature presents a summary of the top 10 submissions in the competition [^6] In this table, the term global forecasting refers to a method where a single model is developed to handle all time series, while cross-validation refers to a technique where the forecasting origin is moved forward by a fixed number of steps, producing multiple forecasts at different points in time..
Although the competition significantly advanced forecasting methodologies, it primarily concentrated on point forecasts, with limited attention to the quantification of forecast uncertainty. Since the competition’s objectives did not explicitly include probabilistic forecasting, criticising these methods for overlooking uncertainty may be unwarranted. Nevertheless, incorporating uncertainty measures is crucial for improving forecast reliability, as decision-makers in the field require dynamic estimates that reflect the uncertainties associated with contraceptive demand over time [43]. The absence of standard approaches for incorporating uncertainty into forecasts highlights a significant gap. While the competition represented a significant step towards the development of advanced forecasting methodologies, it did not fully address the need for probabilistic estimations that can better inform decision making by accounting for forecast variability.
2.2 Human judgment in contraceptive demand forecasting
In reality, uncertainties in contraceptive demand arise from complex patterns of product availability, variable lead times, and the overreliance of donor support [48]. These uncertainties are further complicated by inadequacies in data collection, storage, and sharing practices. Despite the implementation of Logistics Management Information Systems (LMIS), field operatives frequently depend on paper-based forms and Excel spreadsheets for data collection and operational management [20]. This reliance on outdated methods often results in noisy, inaccurate, and incomplete data, thereby complicating the forecasting efforts [11]. Moreover, stockout driven consumption data may not reflect actual demand, further distorting the forecasting process [19].
Forecasting algorithms, while adept at processing high-dimensional data, may struggle to detect these sudden fluctuations and discontinuities in contraceptive demand [33]. Thus, this leaves the expert to use their contextual knowledge to understand the context of the data and incorporate it with the forecasting process [33]. Previous literature also suggests that experts can improve the forecast performance when the external information has not been added to the algorithm-based forecasting method using their inside knowledge and expertise [^7] See Perera et al. [53] for a detailed review on human factors in supply chain forecasting. [22, 17].
In the context of FPSC in developing countries, current algorithm-based forecasting methods possess the capability to manage extensive datasets, including thousands of time series across diverse geographies simultaneously [33]. However, these algorithms may not fully account for external factors and contextual nuances that experts are adept at identifying [20]. This suggests that integrating human expertise with algorithm-based forecasts could yield more reliable and accurate predictions. Empirical evidence also supports the notion that combining human judgment with algorithmic forecasts enhances forecast performance [23, 55]. Furthermore, literature shows that such forecast combinations improve forecasting accuracy compared to using either method in isolation [77]. This leads to the critical question of how best to integrate human expertise with algorithm-based forecasting.
Various approaches exist for integrating human judgment with algorithmic forecasts, forming hybrid intelligence systems that leverage the strengths of both. Brau et al. [13] categorize these methods into five key types:
  • 1.
    Judgmental AdjustmentExperts modify algorithmic forecasts based on contextual insights.
  • 2.
    Quantitative CorrectionSystematic adjustments are applied to human forecasts using statistical techniques.
  • 3.
    Forecast CombinationSeparate judgmental and algorithm-based forecasts are merged into a single forecast.
  • 4.
    Judgment as a Model InputExpert knowledge is incorporated as a predictive variable within the forecasting model.
  • 5.
    Integrative Judgment LearningHuman inputs iteratively refine model predictions through a structured learning process.
Judgmental adjustments have been widely studied in supply chain contexts, particularly in promotional settings. Trapero et al. [71] found that structured human interventions during promotions improved forecasting accuracy. This aligns with findings from Fildes and Petropoulos [24], who argue that human adjustments enhance reliability when bias is minimized.
Quantitative correction methods offer a structured way to adjust forecasts while maintaining statistical integrity. Fildes et al. [23] highlight the importance of systematic corrections to mitigate bias and improve forecast accuracy.
Forecast combination has been explored as a way to merge expert forecasts with algorithmic outputs. Goodwin [28] highlights its effectiveness in capturing both statistical trends and domain knowledge, making it a useful technique when expert input is available alongside data-driven predictions.
Another structured approach to integrating expert knowledge into forecasting models is treating human judgment as a direct input to model-building. Instead of adjusting forecasts post hoc, this method incorporates expert-driven insights as predictive variables, allowing the model to factor in domain knowledge systematically. Arvan et al. [6] suggest that using expert judgment as a model input can improve forecasting accuracy, particularly in cases where statistical models alone struggle to capture contextual nuances. By embedding domain expertise within the forecasting process, this approach ensures that human insights are systematically integrated, rather than relying on manual interventions.
On the other hand, integrative Judgment Learning Baecke et al. [7], Brau et al. [13] treats human adjustments as predictive variables, allowing models to systematically weigh their impact. Similarly, Goodwin and Fildes [29] suggest that hybrid forecasting models should be designed to correct systematic biases in human judgment while maintaining adaptability.
Some studies explore hybrid forecasting methods that do not fit neatly into a single category. These approaches often involve adaptive weighting mechanisms optimization-driven model adjustments, or learning-based frameworks that dynamically integrate human and machine-generated forecasts [54]. While not explicitly categorized under the five methods, these strategies contribute to the broader understanding of how expert judgment and algorithmic forecasts can be effectively combined.
In this context, two key considerations arise when selecting an approach for combining forecasts: 1) healthcare staff at the site level in developing countries often lack the specialized skills and training needed to develop and maintain sophisticated forecasting methods [3], and 2) based on discussions with USAID officials, healthcare staff typically produce only point forecasts, often relying on their judgment, as also highlighted in the literature [2]. Moreover, LaCroix et al. [43] found that although practitioners acknowledge the importance of accounting for uncertainty, there is no widely adopted standard for doing so. Judgmental point forecasts are favored because they align with the practical experience and cognitive abilities of healthcare staff, who may find it difficult to quantify uncertainty without formal training in probabilistic ms. Given these constraints, combining human judgmental forecasts with algorithmic forecasts presents a promising solution. Empirical studies consistently show that combining forecasts, whether judgmental or model-based, generally improves accuracy over relying on individual forecasts alone [58, 77]. However, this introduces a challenge: how to effectively integrate human point forecasts with probabilistic forecasts from algorithms to produce a unified probabilistic forecast. Although the literature on forecast combinations is extensive, little attention has been paid to the integration of point forecasts with probabilistic ones (see Wang et al. [77] for a comprehensive review).
2.3 Literature limitations summary
Despite advancements in contraceptive demand forecasting, several key gaps remain unaddressed. First, the use of probabilistic forecasting in FPSC remains limited. While probabilistic methods have been widely explored in other domains, such as economic forecasting [26, 42], their application in FPSC is largely absent. Most existing studies rely on point forecasts, which do not account for demand uncertainty, making supply chain planning more vulnerable to unexpected fluctuations. Given the inherent unpredictability of contraceptive demand, incorporating probabilistic forecasting is critical for improving stock management and mitigating stockout risks
Second, contraceptive demand forecasting has primarily been conducted at the national or regional level [2, 37], often overlooking demand variability at individual healthcare sites. However, demand patterns vary significantly across locations [38], and without localized forecasts, procurement decisions may fail to align with actual site-level needs. This highlights the need for forecasting approaches that can adapt to site-specific demand while maintaining consistency with broader supply chain strategies.
Third, while forecast combination techniques have been extensively studied [58], their application in FPSC is underexplored. Existing studies typically assess individual forecasting methods without considering how integrating multiple forecasts could improve accuracy, particularly in healthcare supply chains. Additionally, there is limited empirical research comparing different forecasting methods for contraceptive demand estimation, making it difficult to determine the most effective approach in this context.
Finally, although judgmental adjustments are widely used in FPSC [23, 71], there is no structured approach that systematically integrates expert knowledge while ensuring coherence between human forecasts and probabilistic distributions in FPSC settings. In economic forecasting, Entropic Tilting has been used to refine probabilistic distributions based on external point forecasts [42, 46], yet these methods do not explicitly ensure coherence between expert judgment and forecast distributions. Our study addresses this gap by proposing a hybrid forecasting approach that systematically integrates expert forecasts with probabilistic models, ensuring that adjustments remain statistically grounded while incorporating expert insights.
Table 1: Summary of the top 10 models in the USAID intelligent forecasting competition.
Reference Method Metric Forecasting strategy Probabilistic Global forecasting Cross validation
Current study sNAIVE, Moving average, Exponential Smoothing State Space, ARIMA, Syntetos-Boylan approximation, Bayesian structural time series, Multiple Linear Regression, LightGBM, xgBoost, Random Forest, TimeGPT, Chronos, Lag Llama, Demographic, Forecast combination models and Proposed hybrid method MASE, CRPS Recursive multi-step forecasting YES YES YES
Submission 1 An ensemble model of LightGBM and LSTM using weighted average MASE Direct multi-step forecasting NO YES NO
Submission 2 LSTM model MASE Direct multi-step forecasting NO YES NO
Submission 3 A simple ensemble of six LightGBM models MASE Direct multi-step forecasting NO YES NO
Submission 4 LightGBM model MASE Direct multi-step forecasting NO YES NO
Submission 5 LightGBM model MASE Recursive multi-step forecasting NO YES NO
Submission 6 A simple ensemble of six LightGBM models MASE Direct multi-step forecasting NO YES NO
Submission 7 A simple ensemble of LightGBM and LSTM MASE Recursive multi-step forecasting NO YES NO
Submission 8 A simple ensemble of three LightGBM models MASE Direct multi-step forecasting NO YES NO
Submission 9 Hierarchical timeseries model using ARIMA MASE Recursive multi-step forecasting NO NO NO
Submission 10 A simple ensemble of nine LSTM models and Naïve Bayes model MASE Recursive multi-step forecasting NO YES NO
3 Proposed hybrid approach
We propose a CQRA model to generate a combined probabilistic forecast, utilizing both point and probabilistic forecasts. This approach builds upon the CQRA model introduced by Wang et al. [78], which focuses on combining multiple probabilistic forecasts to produce a consolidated forecast distribution. The key concept in our proposed method is to generate quantiles from a given probabilistic forecast and adjust each quantile using weights. These weights are determined by treating the point forecast as the “new reality” and formulating a linear programming (LP) problem that minimizes both the pinball loss [^8] Pinball Loss: A proper scoring rule used to evaluate quantile forecasts, penalizing deviations based on whether the predicted quantile overestimates or underestimates the observed value. It ensures a well-calibrated probabilistic forecast by emphasizing accuracy across different quantile levels. and the absolute error between point forecast and mean of the weighted quantile forecast. The pinball loss is a strictly proper scoring rule used to evaluate quantile forecasts. It measures overall quantile performance by rewarding sharpness and penalizing miscalibration [34].
On the other hand, our proposed CQRA model shares similarities with the quantile combination approach of Trapero et al. [70] in that both methods determine weights by minimizing the tick-loss function. However, a fundamental difference is that while Trapero et al. [70] combine quantile forecasts generated from Kernel Density Estimators (KDE) and GARCH models to optimize safety stock levels, CQRA integrates expert point forecasts with probabilistic forecasts. This allows CQRA to adjust forecast distributions in response to human insights, particularly in settings where domain knowledge complements algorithmic predictions. By explicitly formulating an optimization problem that aligns expert-driven point forecasts with probabilistic distributions, CQRA provides a structured way to incorporate human intuition into data-driven forecasting models, making it well-suited for FPSC and other humanitarian supply chains.
On the other hand, Entropic Tilting has been used as a flexible approach to adjust predictive densities by incorporating external information while maintaining the statistical properties of the original model. This method systematically reweights the probability distribution of forecasts to ensure coherence with expert knowledge or external nowcasts. For example, Krüger et al. [42] applied Entropic Tilting to Bayesian VAR forecasts by incorporating short-term nowcasts, while Metaxoglou and Smith [46] used it to refine option-implied predictive densities for equity returns. These approaches allow for seamless adjustments of predictive distributions while minimizing distortions to the baseline statistical model. While Entropic Tilting provides a suitable method for combining probabilistic and point forecasts, our proposed CQRA model differs in its optimisation framework, which ensures alignment between expert-driven point forecasts and probabilistic forecasts in a constrained manner. Unlike Entropic Tilting, which reweights distributions without necessarily enforcing coherence between expert forecasts and the final probabilistic forecast, CQRA imposes constraints that ensure consistency between the adjusted forecast distribution and expert input. This makes CQRA particularly well-suited for supply chain applications where domain knowledge plays a critical role in refining model-based predictions.
Let the quantile levels be defined as:
$\{q_{1},q_{2},\dots,q_{n}\}\quad\text{where}\quad q_{i}\in[0.01,0.99]$
For a set of weights ${w_{1},w_{2},\dots,w_{n}}$ corresponding to these quantiles, the weighted quantile forecast $\hat{y_{t}}^{(q_{i})}$ for each quantile $q_{i}$ at time $t$ is given by:
$\hat{y}_{t}^{(q_{i})}=w_{i}\cdot\text{ProbForecast}_{t}^{(q_{i})}$
where $\text{ProbForecast}_{t}^{(q_{i})}$ is the probabilistic forecast for quantile $q_{i}$ at time $t$.
The pinball loss for a quantile $q_{i}$ and the point forecast $\text{PointForecast}_{t}$ is defined as:
$L_{q_{i}}(y_{t},\hat{y}_{t}^{(q_{i})})=(y_{t}-\hat{y}_{t}^{(q_{i})})\cdot\left(q_{i}-\mathbf{1}(y_{t}<\hat{y}_{t}^{(q_{i})})\right)$
The weighted mean forecast across all quantiles is calculated as:
$\bar{y}_{t}=\frac{1}{n}\sum_{i=1}^{n}\hat{y}_{t}^{(q_{i})}$
The total loss $L_{t}$ across all quantiles for a single time point $t$ is expressed as:
$L_{t}=\sum_{i=1}^{n}\left|\text{PointForecast}_{t}-\bar{y}_{t}\right|+\sum_{i=1}^{n}\max\left(0,(\text{PointForecast}_{t}-\hat{y}_{t}^{(q_{i})})\cdot\left(q_{i}-\mathbf{1}(\text{PointForecast}_{t}<\hat{y}_{t}^{(q_{i})})\right)\right)$
where; $\sum_{i=1}^{n}\left|\text{PointForecast}_{t}-\bar{y}_{t}\right|$ measures the absolute difference between point forecast and weighted mean quantile forecasts.
The objective is to minimize the total loss $L_{t}$ by optimizing the weights $w_{i}$ across all quantiles:
$\min_{w_{1},w_{2},\dots,w_{n}}\sum_{t=1}^{T}L_{t}$
subject to:
$10\leq w_{i}\leq 1,\quad\sum_{i=1}^{n}w_{i}=$ 1
Once the optimal weights are identified, the final adjusted quantile forecast $\tilde{y}_{t}^{(q_{i})}$ for each quantile $q_{i}$ is:
$\tilde{y}_{t}^{(q_{i})}=w_{i}^{*}\cdot\text{ProbForecast}_{t}^{(q_{i})}$
However, probabilistic forecasts may have a more dominant influence in this approach since the weights $w_{i}$ are restricted to the range $[0,1]$. This implies that point forecasts are expected to align closely with the forecast distribution. When the probabilistic forecasts are reliable and the point predictions do not significantly deviate from the mean of the probabilistic forecast, this approach is recommended. However, recognizing that this is often not the case in practice, we relax the sum-to-one constraint on the combination weights to provide greater flexibility in incorporating expert point forecasts and accounting for bias. A similar relaxation has been explored in prior research (e.g., [30]), which suggests that imposing weight constraints can sometimes lead to suboptimal forecasts by failing to account for systematic biases. Furthermore, Trapero et al. [70] also show that an unconstrained approach to optimizing quantile forecast combinations can yield improvements, particularly in applications requiring adaptive weight adjustments. By allowing weights to exceed unity and introducing a bias factor, our modified CQRA approach accommodates situations where expert forecasts systematically deviate from model-based predictions, enhancing the adaptability and robustness of the combined forecast.
Modified Approach:
  1. 1.
    We update the weighted quantile forecast $\hat{y}_{t}^{(q_{i})}$ as follows:
$\hat{y}_{t}^{(q_{i})}=w_{i}\cdot\text{ProbForecast}_{t}^{(q_{i})}+b_{t}$
where $b_{t}$ is the bias factor at time $t$ and it is calculated as the parameter optimized alongside the weights.
  1. 2.
    We remove the normalization constraint $\sum_{i=1}^{n}w_{i}=1$ and increase the upper bound of $w_{i}$ to 5:
$0\leq w_{i}\leq 5$
Our sensitivity analysis showed that increasing the upper bound beyond 5 did not significantly improve method performance, making 5 an optimal choice for balancing flexibility and control.
  1. 3.
    After optimization, we apply an adjustment factor to ensure that the mean of the forecast distribution aligns with the point forecast:
$\tilde{y}_{t}^{(q_{i})}=w_{i}^{*}\cdot\text{ProbForecast}_{t}^{(q_{i})}\cdot adj_{t}$
The adjustment factor $adj_{t}$ is defined as:
$adj_{t}=\frac{\text{PointForecast}_{t}}{\bar{y}_{t}}$
These adjustments ensure that the combined probabilistic forecast aligns with the central tendency of the point forecast while still capturing the uncertainty in the prediction. This approach is particularly useful when the probabilistic forecast does not include external variables that cause significant deviations in demand.
After generating the weighted quantile forecast $\tilde{y}_{t}^{(q_{i})}$ for each quantile $q_{i}$, we create a smooth forecast distribution by linearly interpolating between the quantile levels:
$\tilde{y}_{t}(x_{j})=\tilde{y}_{t}^{(q_{i})}+\left(\frac{x_{j}-q_{i}}{q_{i+1}-q_{i}}\right)\cdot\left(\tilde{y}_{t}^{(q_{i+1})}-\tilde{y}_{t}^{(q_{i})}\right)$
where $q_{i}\leq x_{j}<q_{i+1}$ and $x_{j}$ are the interpolation points.
The final interpolated forecast distribution $\tilde{Y}_{t}$ for all interpolation points $x_{j}$ is:
$\tilde{Y}_{t}=\{\tilde{y}_{t}(x_{1}),\tilde{y}_{t}(x_{2}),\dots,\tilde{y}_{t}(x_{m})\}$
Remark 1: When providing point forecasts to the method, they should first be combined with the mean forecasts from the probabilistic forecast using a simple averaging method. This combined point forecast will serve as the new central tendency (e.g., mean or median) for the overall forecast. Based on this combined central tendency, optimal weights will then be determined to enhance the accuracy and balance of the forecast by integrating both the point and probabilistic perspectives effectively.
Remark 2: We refer to the first proposed combination method as the Hybrid Weighted Average, and the revised version of the combination method is termed the Hybrid Bias Adjustment.
Underlying assumptions of the method are;
  1. 1.
    The probabilistic forecasts are well-calibrated.
  2. 2.
    The point forecast accurately represents the central tendency of the future distribution.
  3. 3.
    The weights used for each quantile are restricted to non-negative values, ensuring that the final forecast distribution remains in the range of the original probabilistic forecasts.
  4. 4.
    A linear combination of quantile forecasts, point forecasts, and bias adjustment is sufficient to represent the true forecast distribution.
  5. 5.
    Linear interpolation between quantile levels accurately reflects the true underlying distribution.
4 Experiment setup
4.1 Data collection and preprocessing
The data used in our study were extracted from the LMIS of Cote d’Ivoire. The dataset encompassed 156 sites distributed across 81 districts in 20 regions, covering a span of 11 contraceptive products across 7 product categories. These categories included female and male condoms, emergency contraceptives, oral contraceptives, injectables, implants, and IUDs. The dataset spanned from January 2016 to December 2019 at a monthly granularity containing 1454 time series. Figure 1 shows the location of each site in Côte d’Ivoire by site type, illustrating that the sites are distributed throughout the country.
Our initial exploration indicated that there were no duplicate values; however, some missing values were present in the time series. Additionally, a few time series contained stockout cases. Since our study does not focus on handling stockouts in the forecasting process, we removed the series with stockouts and missing values, as we could not determine the reasons behind those missing values. This filtering resulted in a final dataset of 1,360 time series.
In our study, we focus on stock distributed [^9] We assume that stock distributed serves as a reasonable proxy for consumption data, as we eliminated stockout cases due to limited access to direct consumption data from users. as the target variable at the site level for various contraceptive products.
Figure 1: Contraceptive stock distribution in Côte d’Ivoire by healthcare site location. The size of the circles represents the quantity of stock distributed.
4.2 Data exploration
First, we examined the time plots of the data at various aggregation levels to observe the time series features such as trend, seasonality, and noise. As illustrated in Figure 2, higher aggregation levels reveal clearer seasonal patterns and trends, while lower aggregation levels exhibit increased volatility. Additionally, the plots highlight notable differences in stock distribution across locations and products, suggesting the presence of distinct patterns associated with each.
At the lowest aggregation level, depicted in Figure 3, which focuses on product distribution at individual site level, demand patterns become more heterogeneous, comprising a mix of smooth, erratic, lumpy, and intermittent demand types. Unlike the aggregate levels, where trends and seasonality are more apparent, these patterns are not easily discernible at the site level, further complicating the forecasting process.
Figure 2: Time series of contraceptive product stock distributed (Jan 2016 – Dec 2019) at various levels. The x-axis represents the month, while the y-axis indicates the number of units distributed. The panels display data from the entire country (top panel), with breakdowns by region, district, site, and product code. The bottom panel shows the number of units distributed in selected sites for specific products. To ensure clarity and prevent overplotting, only five time series are displayed for each aggregate level. These series were selected randomly and are characteristic of the patterns encountered at the respective aggregation levels.
Figure 3: Time series of contraceptive product stock distributed in selected sites for specific products(Jan 2016 – Dec 2019). To ensure clarity and prevent overplotting, only five of the products are displayed. These series were selected randomly and represent characteristic patterns at this level.
Next, we examined the time series data of the products at the site level to gain a clearer understanding of trend and seasonality patterns, as our primary focus is on forecasting each product at the site level. However, due to the large number of time series, it was not visually feasible to plot all individual series together to simultaneously inspect trends and seasonality. Therefore, we employed the Seasonal and Trend Decomposition using Loess (STL) method [16] to extract key features from all 1,360 time series.
As shown in Figure 4, the strength of trend and seasonality for each time series is represented on a scale from 0 to 1, where 0 indicates low strength and 1 indicates high strength. The majority of the time series exhibited moderate levels of both trend and seasonality. However, even within the same product code, we observed variations in trend and seasonality patterns, which posed challenges for the forecasting process. Consequently, we considered a range of forecasting approaches, including time series, Bayesian, ML, and foundational time series methods, to determine which could most effectively handle the diverse patterns within the time series.
Figure 4: Trend strength and seasonality in the time series of stock distribution. Each point in the scatter plot represents one of the 1,360 time series analysed, with trend and seasonality strengths measured on a scale from 0 to 1 (0 indicating weak and 1 indicating strong).
4.3 Forecasting setup
Our forecast setup began with data collection and preparation of a tidy dataset for the forecasting process. Following this, we carried out feature engineering. As outlined by Kolassa et al. [41], incorporating lag predictors and rolling window statistics is beneficial for improving forecasting methods. In addition to these, we also integrated categorical and date features into the forecasting process. To ensure we selected the most relevant variables, a feature importance analysis was conducted to identify the best predictors for the forecasting methods.
In the USAID forecasting competition, the planning horizon was set to 3 months. However, instead of using fixed training and test sets as in the competition, we adopted the time series cross-validation approach to create the training and test sets [34]. Unlike the fixed approach, where the same training and test sets are used for evaluation, time series cross-validation moves the forecasting origin forward by a fixed number of steps, producing multiple forecasts at different points in time. This allows for the calculation of multi-step errors, giving a more robust view of how methods perform across various demand scenarios, such as periods of high and low demand [66].
In our cross-validation setup, we define the training period as all available data up to September 2019, ensuring sufficient historical observations for model learning. The testing period consists of rolling evaluation windows, where each test set spans a 3-step-ahead forecasting horizon to align with the competition requirements. At each iteration, the training set expands while the test set moves forward by one step, maintaining a consistent evaluation structure. We limited the number of rolling origins to 3 per series due to computational constraints, but this still provided us with meaningful insights into method performance over time. For forecasting, we employed recursive multi-step forecasting and we generated 1000 future paths per a series. All method development and hyper-parameter tuning were conducted using only the training data to ensure that the evaluation remained unbiased and the methods were properly validated.
4.3.1 Probabilistic forecasting using bootstraping
To express the uncertainty of our forecasting methods’ estimates, we utilize probability distributions of potential future values. Several methods are available to estimate prediction intervals, including analytical prediction intervals, bootstrapping, quantile regression, Bayesian modeling (using MCMC sampling), and conformal prediction. In our study, we employ the bootstrapping method to estimate these intervals [26].
Given that our study employs multiple forecasting methods, bootstrapping provides a unified framework that can be applied consistently without requiring additional parametric assumptions. Moreover, it allows us to approximate the empirical distribution of forecast errors without imposing strict distributional constraints, making it particularly suitable for datasets with varying demand patterns in the FPSC.
To account for uncertainty in predictions, we assume that future errors will resemble past errors. We define the error as the difference between the actual value and the forecasted value:
$e_{t}=y_{t}-\hat{y}_{t}$
where; $e_{t}$ is the error at time $t$, $y_{t}$ is the actual value and $\hat{y}_{t}$ is the forecast value at time $t$.
We simulate different future predictions by sampling from the collection of past errors and adding these to the forecast estimates. Each bootstrap iteration produces a different potential future path. By repeating this process, we generate a distribution of possible outcomes. Based on a chosen significance level, prediction intervals can then be constructed from this distribution [34].
To implement the bootstrapping process, we use the fable package in R for time series methods and the skforecast package for ML methods [4]. However, for the Bayesian methods, this process is not necessary as it inherently provides probabilistic forecasts as part of its output. Similarly, foundational time series methods also deliver probabilistic forecasts directly.
4.3.2 Forecast combination
Forecast combination is a promising approach to enhance forecasting performance by aggregating multiple forecasts generated using different methods for a specific time series. This technique eliminates the need to select a single “best” forecasting method, thus leveraging the strengths of various methods [77]. Known as either forecast combination or forecast ensemble across different fields, this method has been widely used and extensively studied [27]. The literature provides substantial evidence that forecast combinations consistently outperform individual forecasts [58], primarily by mitigating uncertainties arising from data variability, parameter estimation, and method selection [77].
Forecast combination methods can range from linear combinations, nonlinear combinations, and time-varying weights, to more sophisticated approaches like cross-learning, correlations among forecasts, or Bayesian techniques [77]. Among these, the most widely adopted approach is the linear combination with equal weights [27]. This method is not only straightforward to implement and interpret but also provides robust and improved forecasting performance [58, 27, 69]. Consequently, in our study, we applied a linear combination approach with equal weights to generate combined forecasts.
4.4 Forecasting methods
In our study, we employed a range of forecasting methods to address the volatile nature of the time series data. For time series methods, we used sNAIVE, Moving Average (MA), Exponential Smoothing State Space (ETS), ARIMA, and Syntetos-Boylan approximation (SBA). As Bayesian methods, we implemented the Bayesian Structural Time Series (BSTS) method with regressors. For ML methods, we applied Multiple Linear Regression (MLR), Random Forest (RF), LightGBM (LGBM), and XGBoost (XGB). These methods were selected due to their popularity, efficiency, and ease of implementation within the forecasting domain [44]. Furthermore, for the ML methods, we developed each as a global method, where a single method was trained to produce forecasts for all time series simultaneously [9].
Additionally, we explored foundational time series approaches such as TimeGPT [25], Chronos [5] and Lag Llama [59], which are gaining attention due to advancements in large language models (LLMs) and also capable of producing probabilistic forecasts. These methods offer zero-shot forecasting capabilities, meaning they have been pre-trained on vast amounts of time series data and can be applied to new time series without the need for retraining or fine-tuning parameters [15]. This feature significantly reduces the steps typically required in the forecasting process, such as data preparation, model training, and model selection [25]. However, these methodologies have yet to be tested within the FPSC context.
To offer a more comprehensive comparison of forecasting methods in contraceptive demand forecasting, we incorporated a demographic forecasting method. This method uses demographic data such as population size, age distribution, and other family planning indicators to estimate future contraceptive demand [2]. Given that we did not have access to the final forecasts generated at the site level by demand planners, we assumed that the demographic-based method serves as a proxy for expert-driven forecasts. This assumption is grounded in the fact that experts typically leverage their domain knowledge when determining key family planning indicators.
4.4.1 Time series methods
sNAIVE: This method is a simple forecasting approach where forecasts are generated using the most recent observation from the corresponding period of the previous cycle. This method is often used as a benchmark in time series forecasting [34] and which can be shown as;
$\hat{y}_{t+h}=y_{t+h-s}$
Where $\hat{y}_{t+h}$ is the forecast for $t+h$, and $s$ is the seasonality period. We implemented this method using the SNAIVE() function in the fable package in R [51].
MA: MA method is a simple forecasting approach that generates predictions by averaging a fixed number of the most recent observations. This method helps to smooth out short-term fluctuations while emphasizing longer-term trends in the data. The method assumes that future values can be reasonably estimated based on the mean of past values over a specified window [34]. MA method can be represented as:
$\hat{y}_{t+1}=\frac{1}{n}\sum_{i=0}^{n-1}y_{t-i}$
where $\hat{y}{t+1}$ is the forecast for the next time period, $n$ is the number of past observations (the window size), and $y{t-i}$ are the actual values from previous periods.Due to the simplicity nature of this method, it is often used in the FPSC context. We implemented this method using the MEAN() function in the fable package in R [51]
ETS: ETS model accommodates trends, seasonality, and error terms within time series data through various approaches, such as additive, multiplicative, or mixed models within a state-space framework. The model updates these components dynamically over time using recursive equations. The ETS model is capable of handling diverse time series patterns, including trends and seasonal fluctuations [34]. Given the large number of series in our dataset, we utilized the automated ETS model, which selects the optimal model based on Akaike’s Information Criterion (AIC) for each time series. We used the ETS() function in the fable package in R [51] to implement this model.
ARIMA: ARIMA model forecasts based on trends, autocorrelation, and noise within time series data. It is also flexible and can handle both non-seasonal and seasonal data by incorporating seasonal components. ARIMA parameters (p,d,q) denote the orders of the auto-regressive (AR) component, differencing, and moving average (MA) component, respectively. ARIMA is particularly effective for data with a pronounced temporal structure [34]. Like with the ETS model, we employed an automated approach to fit ARIMA models for each time series using the ARIMA() function in the fable package in R, which selects the best model using similar criteria [51].
SBA: Since some time series exhibit an intermittent demand nature, we also employed the SBA method in our study, an enhancement of Croston’s original method from 1972 [67]. The SBA approach methods intermittent demand as a binomial process by separately estimating the demand intervals and the demand sizes when they occur. This method applies a correction factor to reduce the inherent positive bias of the original Croston method, making the forecasts more accurate. We implemented this method using the CROSTON(type = ‘sba’) function in the fable package in R [51].
4.4.2 Bayesian methods
BSTS: The BSTS method used in our study combines a local linear trend and a seasonal component, incorporating additional covariates to fit the observed data [40]. The local linear trend is a time-varying method that captures the evolving pattern of the time series over time. It consists of a level and a slope, both of which are allowed to change dynamically. The state equations for this are:
$\mu_{t}=\mu_{t-1}+\beta_{t-1}+\eta_{t},\quad\eta_{t}\sim N(0,\sigma_{\eta}^{2})$
$\beta_{t}=\beta_{t-1}+\zeta_{t},\quad\zeta_{t}\sim N(0,\sigma_{\zeta}^{2})$
where; $\mu_{t}$ is the level at $t$, $\beta_{t}$ is the slope at $t$, $\eta_{t}$ and $\zeta_{t}$ are normally distributed errors with variances $\sigma_{\eta}^{2}$ and $\sigma_{\zeta}^{2}$ respectively.
The seasonal component captures regular patterns that repeat over a fixed period and it is modeled as:
$S_{t}=-\sum_{j=1}^{m-1}S_{t-j}+\omega_{t},\quad\omega_{t}\sim N(0,\sigma_{\omega}^{2})$
where; $S_{t}$ is the seasonal effect at time $t$, $m$ is the number of seasons (in our case, $m$ = 12 for monthly data) and $w_{t}$ is the normally distributed error with variance $\sigma_{\omega}^{2}$.
The observed data (i.e., target variable) is modeled as a linear combination of the local linear trend, seasonal component, and additional regressors. This is represented by the observation equation:;
$y_{t}=\mu_{t}+S_{t}+X_{t}\beta+\epsilon_{t},\quad\epsilon_{t}\sim N(0,\sigma_{\epsilon}^{2})$
where; $X_{t}$ are the regressors, $\beta$ are the corresponding coefficients and $\epsilon_{t}$ is the noise.
Posterior distributions for parameters are estimated using Markov Chain Monte Carlo (MCMC) methods [40]. The predicted future values are obtained by simulating from these posterior distributions and thus quantifying the uncertainty given the Bayesian nature of the method [45].
4.4.3 ML methods
MLR: MLR methods establish linear relationships between the target variable and multiple predictor variables. The method estimates coefficients for each predictor variable by minimizing the residual sum of squares between observed and predicted values. These methods are particularly useful when demand is influenced by various factors [34]. In our study, we implemented this method using the LinearRegression() function from the sklearn package in Python [52].
RF: RF is an ensemble learning method that constructs a collection of decision trees, each trained on a bootstrap sample of the original data. The predictions of these trees are aggregated to produce the final forecast [14]. We used the RandomForestRegressor() function from the sklearn package in Python [52] to implement the RF method.
Gradient Boosted Regression Trees (LGBM and XGB): These methods are known for their efficiency and ease of implementation [44]. These methods use an ensemble of decision trees, where each new tree is added to correct the residuals of the previous trees in an iterative manner [35]. Unlike Random Forest, which builds trees independently, gradient boosting methods focus on improving method performance iteratively. In our study, we selected LightGBM (LGBM) and XGBoost (XGB) for their ability to handle multiple predictor variables in various forms (binary, categorical, and numeric) and their effectiveness in providing reliable and accurate forecasts [44]. We used the LGBMRegressor() function from the LightGBM package in Python [47] and the XGBRegressor() function from the XGBoost package in Python [21]. Hyperparameter tuning for both LGBM and XGB was performed using grid search, with the Poisson distribution chosen as the objective function due to the count nature of the target variable.
4.4.4 Foundational time series methods
TimeGPT: TimeGPT is the first pre-trained foundational method specifically designed for time series forecasting, developed by Nixtla [25]. It employs a transformer-based architecture with an encoder-decoder setup, but unlike other methods, it is not derived from existing large language methods (LLMs); rather, it is purpose-built to handle time series data. TimeGPT was trained on over 100 billion data points, encompassing publicly available time series from a variety of domains, including retail, healthcare, transportation, demographics, energy, banking, and web traffic. Due to the diversity of these data sources and the range of temporal patterns they exhibit, TimeGPT can effectively handle a wide variety of time series characteristics. Additionally, the method can incorporate external regressors into the forecasting process and is capable of producing quantile forecasts, allowing for robust uncertainty estimation (see Ansari et al. [5] for a detailed overview).
Chronos: Chronos is a univariate probabilistic foundational time series method developed by Amazon [5]. Like TimeGPT, it is based on a transformer architecture in an encoder-decoder configuration, but it trains an existing LLM architecture using tokenized time series via cross-entropy loss. Chronos was pre-trained on a large publicly available time series dataset, as well as on simulated data generated through Gaussian processes. The method was trained on 28 datasets, comprising approximately 84 billion observations. Chronos is based on the T5 family of methods, offering different versions with parameter sizes ranging from 20 million to 710 million. The four pre-trained methods available for forecasting are: 1) Mini (20 million), 2) Small (46 million), 3) Base (200 million), and 4) Large (710 million) (see Ansari et al. [5] for a detailed overview). In our study, we employed the Base Chronos T5 method for its balance between performance and computational efficiency.
Lag Llama: Lag Llama is another univariate probabilistic foundational time series method, which is based on the LLaMA architecture and utilizes a decoder-only structure [59]. The method tokenizes time series data using lags as covariates and applies z-normalization at the window level. This approach focuses on learning time series behavior from past observations. Lag Llama was trained on 27 publicly available time series datasets across six domains: nature, transportation, energy, economics, cloud operations, and air quality. With 25 million parameters, this method is designed to handle diverse time series frequencies and features, making it suitable for a wide range of forecasting tasks (see Rasul et al. [59] for a detailed overview).
4.4.5 Demographic forecasting method
In the FPSC context, the demographic forecasting method is employed to estimate contraceptive needs for a given population based on a set of family planning indicators during the forecast period. This method is formulated as a combination of these indicators and population dynamics, as represented in the following equation [2]:
$y_{i,t}=\left(\sum_{j=15}^{50}\left(mCPR_{t,j}\times WomenPopulation_{t,j}\right)\right)\times MethodMix_{t,i}\times CYP_{t,i}\times BrandMix_{t,i}\times SourceShare_{t}$
Where $i$ represents the product, $j$ is the age group, $mCPR$ is the modern contraceptive prevalence rate, and $CYP$ refers to couple-years of protection.
$WomenPopulation_{t,j}$ denotes the total population of women in a selected location, typically within the age range of 15-49 years, which is the standard range used in census data or demographic health surveys. For our study, we sourced this population data from WorldPop [79] and mapped it to each healthcare site based on the latitude and longitude coordinates of those sites.
$mCPR$ stands for the percentage of women of reproductive age using modern contraceptives, with data collected from the PMA Data Lab [56].
$MethodMix$ represents the share of different contraceptive methods being used, including injectables, IUDs, implants, pills, and condoms. This data is also obtained from the PMA Data Lab [56].
$CYP$ is a metric estimating the protection from pregnancy provided by a contraceptive method over one year. For example, an implant can cover 3.8 years, so CYP adjusts for such longer-acting methods. We collected this data from USAID [75].
$BrandMix$ reflects the brand share percentage within each contraceptive method. This was calculated using historical data.
$SourceShare$ refers to where women of reproductive age, using a specific method and brand, obtain their products. This mix typically includes public, private, NGO/SMO (social marketing organizations), and other small providers. Data was gathered through discussions with USAID officials.
This equation provides $y_{i,t}$, which is the total annual point estimate of contraceptives required for product $i$ at time $t$. It is typically used at the national level on an annual basis to inform procurement decisions [2].
However, as our study focuses on monthly estimates at the healthcare site level, we revised the equation by introducing a weighting factor, $w_{t}$, to distribute the annual estimates across months. The revised equation is as follows:
$y_{i,t,s}=\left(\sum_{j=15}^{50}\left(mCPR_{t,j}\times WomenPopulation_{t,j,s}\right)\right)\times MethodMix_{t,i}\times CYP_{t,i}\times BrandMix_{t,i}\times\ SourceShare_{t}\times w_{t}$
Where $w_{t}$ represents the monthly weight, $s$ is the healthcare site, and $y_{i,t,s}$ is the monthly point forecast for product $i$ at healthcare site $s$.
4.4.6 Overview of candidate methods
In our study, we developed 20 candidate methods by experimenting with different combinations of predictors and by combining various forecasting methods. For the MA method, we opted to use a three-month averaging period, aligning with the current practice at the site level in Côte d’Ivoire. Additionally, we developed two model combinations using equal-weight linear averaging: a combined statistical model and a combined ML model.
To create hybrid probabilistic methods, we combined point forecasts from the demographic method with the combined ML method, resulting in a hybrid combined method. This hybrid method synthesizes insights from both the demographic point forecast and the probabilistic algorithm-based forecast, aiming to capture expert knowledge alongside data-driven characteristics of machine learning methods. This integration is intended to enhance forecast accuracy by leveraging the strengths of both approaches.
We developed two variations of the hybrid probabilistic method based on our proposed methods. A detailed overview of all 20 candidate methods is provided in Table LABEL:tbl-models. We also explored several other approaches to develop different forecast method variations. These included using demographic indicators as predictors, different method combinations and applying hierarchical forecasting reconciliation to combine demographic-based forecasts with algorithm-based forecasting methods. However, we decided not to include the results of these methods, as they did not improve the performance significantly.
Table 2: Proposed candidate methods in our study
Type Method Predictor variables Remarks Probabilistic Forecasts
Time series sNaive Historical stock distributed data - Yes
Moving average Historical stock distributed data - Yes
ETS Historical stock distributed data - Yes
ARIMA Historical stock distributed data - Yes
Croston-SBA Historical stock distributed data - No
Bayesian BSTS reg Historical stock distributed data, lag values (for 1,2,3,4), lag rolling mean, 4 period rolling max, 4 period rolling zero percentage, Month and year, region, district, site type, site code, product type and product code - Yes
BSTS demo Historical stock distributed data, women population at each site, mCPR, method mix,CYP, brand mix, source share - Yes
ML MLR Historical stock distributed data, lag values (for 1,2,3,4), lag rolling mean, 4 period rolling max, 4 period rolling zero percentage, Month and year, region, district, site type, site code, product type and product code - Yes
RF - Yes
LGBM - Yes
XGB - Yes
Demographic Demographic Women population at each site, mCPR, method mix,CYP, brand mix, source share, weight for each month - No
Foundational TimeGPT Historical stock distributed data - Yes
TimeGPT reg Historical stock distributed data, lag values (for 1,2,3,4), lag rolling mean, 4 period rolling max, 4 period rolling zero percentage, Month and year, region, district, site type, site code, product type and product code - Yes
Chronos Historical stock distributed data - Yes
Lag Llama Historical stock distributed data - Yes
Combination Statistical combined - Model combination using sNAVIE, MA, ETS, ARIMA. We didn’t use Croston as it only produces point forecats. Yes
ML combined - Model combination using RF, LGBM, XGB. We didn’t use MLR because it significantly reduces combined forecast performance. Yes
Hybrid Hybrid weighted average - Combination between demographic mdethod and ML combination using the weighted average approach. Yes
Hybrid bias adjustment - Combination between demographic method and ML combination using the weighted average bias approach. Yes
4.5 Performance evaluation
To assess the performance of our forecasting methods, we used both point forecast and probabilistic forecast evaluation metrics. We evaluated point forecasts using the MASE. MASE was chosen for two primary reasons. First, it was the official evaluation metric used in the USAID Intelligent Forecasting Competition, allowing us to directly compare our model performance with previous benchmark results. Second, MASE is a scale-independent metric that provides robustness, and stability [41].
The MASE formula is:
$\text{MASE}=\text{mean}(|q_{t}|),$
where
$q_{t}=\frac{e_{t}}{\displaystyle\frac{1}{n-m}\sum_{t=m+1}^{n}|y_{t}-y_{t-m}|},$
Here, $e_{t}$ is the point forecast error for forecast period $t$, $m=12$ (to account for seasonality), $y_{t}$ is the observed value, and $n$ is the number of observations in the training set. The denominator is the mean absolute error of the seasonal naive method over the training sample, ensuring the error is properly scaled. Smaller MASE values indicate more accurate forecasts, and since it was the metric used in the USAID competition, it allows us to compare our results with the competition submissions.
To evaluate the accuracy of probabilistic forecasts, we employed the CRPS, a widely used metric in probabilistic forecasting that assesses the sharpness and calibration of the forecast distribution.
The CRPS is given by:
$\text{CRPS}=\text{mean}(p_{j}),$
where
[…]
where $G_{t}(x)$ is the forecasted probability distribution function for the period $t$, and $F_{t}(x)$ is the true probability distribution function for the same period.
CRPS is beneficial to our study as it measures the overall performance of the forecast distribution by rewarding sharpness and penalizing miscalibration [26]. Calibration measures how well predicted probabilities match the true observations, while sharpness focuses on the concentration of the forecast distributions [77]. Thus, CRPS provides a single score by evaluating both calibration and sharpness, making it easy to evaluate the performance of forecasting methods. In this formula, $G_{t}(x)$ is the forecasted cumulative distribution function (CDF) for time $t$ and $F_{t}(x)$ is the true CDF for the same time. The CRPS evaluates the difference between the predicted and actual probability distributions, with lower values indicating better performance [58]. It combines aspects of both calibration (the alignment of predicted probabilities with actual outcomes) and sharpness (the concentration of the forecast distribution), making it a comprehensive measure of forecast quality [77].
While CRPS provides a comprehensive evaluation of the entire predictive distribution, there are cases where accuracy at specific quantiles is of particular interest. For example, in inventory management of family planning health commodities, higher quantiles (e.g., the 95th percentile) might be important to ensure efficient stock management and maintain a high service level. By accurately capturing demand at these upper quantiles, supply chain planners can better mitigate stockouts and ensure the consistent availability of essential health supplies. In such scenarios, the quantile score (or pinball score) becomes a more appropriate metric, as it directly evaluates forecast accuracy at the chosen quantile, ensuring a more targeted assessment of predictive performance.
5 Analysis and results
First, we evaluate the overall point forecast performance of the forecasting methods, including the proposed method, using the MASE. Additionally, we compare the overall performance of our methods against the top 10 submissions from the USAID competition. Second, we assess the overall performance of the probabilistic forecasts of our methods using the CRPS. After completing these evaluations, we conduct a Nemenyi test at the 5% significance level to determine any significant differences in performance between the methods.
Next, we evaluate both the point and probabilistic forecast performances across forecast horizons, providing a clearer picture of multi-step errors in the methods. Following this, we compare the forecast performances in relation to computational time, highlighting the trade-offs between accuracy and efficiency.
5.1 Overal performance evaluation of point and probabilistic forecasts [^10] Overall performance refers to the mean and median forecast performance of methods calculated on the test sets at forecast horizons $h=1,2,3$ months, with time series cross-validation applied to the target variable.
The overall point forecast performance of each method is presented in Table 3, showing both mean and median MASE values, and ordered by mean MASE. The table clearly indicates that all time series methods underperform compared to ML methods. In fact, the top five methods are ML-based. The top-performing method is the RF method, with the lowest mean MASE of 0.743. Notably, the Hybrid Weighted Average method is the second-best performer, with a mean MASE of 0.775.
However, the Hybrid Bias Adjustment method performs significantly worse compared to all other methods, except for the Demographic method, which shows the poorest performance among all methods. Interestingly, the SBA method outperforms all other time series methods, but neither the Statistical Combined method nor the ML Combined method surpass other methods within their respective categories as initially expected. Nevertheless, it is notable that both combined methods improve their performance compared to the lowest-performing methods within their category.
On the other hand, the BSTS method shows improved performance when it incorporates time series-based predictors (e.g., lags, rolling statistics), categorical features (e.g., region, district), and date features, compared to when it uses demographic-based predictors (e.g., women population, mCPR, women age group). Among the foundational methods, TimeGPT with regressors outperforms all other foundational methods, whereas without regressors, the performance of Chronos and TimeGPT is quite similar. However, the performance of Lag Llama differs notably from both Chronos and TimeGPT. Finally, regarding the competition submissions, none of them outperform the top five methods in our analysis.
Table 3: Overall point forecast accuracy in mean MASE and median MASE (CS refers to competition submission).
Method Mean MASE Median MASE
RF 0.743 0.376
Hybrid weighted averaging 0.775 0.426
LGBM 0.833 0.426
ML combined 0.847 0.460
XGB 0.859 0.433
CS 01 0.990 0.789
CS 02 0.995 0.798
CS 03 0.998 0.779
CS 04 1.004 0.790
CS 05 1.014 0.815
CS 06 1.035 0.785
CS 07 1.043 0.823
CS 08 1.051 0.819
CS 09 1.088 0.844
CS 10 1.103 0.861
TimeGPT reg 1.258 0.623
MLR 1.269 0.632
TimeGPT 1.292 0.669
Chronos 1.305 0.641
BSTS reg 1.327 0.732
SBA 1.331 0.689
Moving average 1.373 0.694
Statistical combined 1.378 0.731
ETS 1.379 0.683
ARIMA 1.386 0.689
Lag Llama 1.483 0.777
BSTS demo 1.521 0.859
sNaive 1.603 0.924
Hybrid bias adjustment 4.360 0.800
Demographic 16.072 1.847
However, we cannot draw concrete conclusions about the point forecast performance of methods solely based on mean MASE values. Therefore, we also conducted the Nemenyi test at the 5% significance level on MASE values for the forecasting methods. This test allowed us to calculate the average ranks of the forecasting methods and assess whether their performances are significantly different from one another. Figure 5 shows the results of the Nemenyi test.
In brief, if there is no overlap in the confidence intervals between two methods, it indicates that their performances are significantly different. The grey area represents the 95% confidence interval for the top-ranking method. Methods whose intervals do not overlap with this grey area are considered significantly underperforming compared to the best-performing method, and vice versa.
Figure 5 demonstrates that the RF method is the best-performing method confirming our previous finding, and there is no significant difference between the top three ranked methods, which include our proposed Hybrid Weighted Average method and the LGBM method. It is noteworthy that the average rank of the Hybrid Bias Adjustment method has improved, suggesting that it may perform adequately across a majority of the time series. Additionally, it is significant that the TimeGPT with Regressors method outperforms all other foundational time series methods, which were trained as univariate methods.
Figure 5: Average ranks of forecasting methods with 95% confidence intervals based on the Nemenyi test for MASE values. Lower ranks indicate better forecast performance.
Next, we turn our attention to evaluating the performance of probabilistic forecasts. Table LABEL:tbl-crps presents the overall performance evaluations of probabilistic forecasts using both the mean and median CRPS values, ordered by mean CRPS. The proposed Hybrid Weighted Averaging method is the top performer, with a mean CRPS of 9.868. The RF method ranks second, with a mean CRPS of 9.997. As in the point forecast analysis, all the top five methods are ML based, and the time series methods generally underperform in comparison.
In the BSTS method, we again observe improved performance when time series-based, categorical, and date features are included as regressors. The Statistical Combined and ML Combined methods show performance similar to what was seen in the point forecast analysis.
Notably, Chronos performs better than all time series, Bayesian, and other foundational methods. Moreover, ETS outperforms all time series methods but shows poor performance compared to ML based methods. Lastly, the Hybrid Bias Adjustment method delivers the worst performance among all forecasting methods, reinforcing the trend observed in the point forecast evaluation.
Table 4: Overall probabilistic forecast accuracy in mean CRPS and median CRPS.
Method Mean CRPS Median CRPS
Hybrid weighted averaging 9.868 3.083
RF 9.997 2.754
LGBM 10.131 3.067
ML combined 10.286 3.377
XGB 10.560 3.164
MLR 12.611 4.512
Chronos 15.018 4.698
BSTS reg 15.342 5.275
ETS 15.397 5.632
TimeGPT reg 15.635 4.783
Moving average 15.701 5.480
ARIMA 15.703 5.602
TimeGPT 15.831 5.275
Lag Llama 15.840 5.919
Statistical combined 16.045 5.671
BSTS demo 17.064 6.526
sNaive 17.511 7.119
Hybrid bias adjustment 29.062 6.447
Similar to the point forecast analysis, Figure 6 demonstrates that the top three ranked methods are not significantly different, with RF as the top-ranked method, although the proposed Hybrid Weighted Average method has the lowest mean CRPS. This may indicate that RF performs comparably in minimizing the loss function across series, while the Hybrid Weighted Average method may prioritize stable time series without significant deviations (see Section 3). Additionally, the top three ranked methods significantly outperform all other forecasting methods in terms of probabilistic forecasting.
Noticeably, the Hybrid Bias Adjustment method shows a significant improvement in its average rank, ranking seventh, right after the ML and Hybrid Weighted Averaging methods. The Chronos method is also ranked higher than the time series, BSTS, and other foundational methods. Time series methods remain clustered in the lower rank range, while the BSTS reg method shows an improvement in rank compared to the BSTS demo method. Among the foundational methods, Lag Llama has the lowest rank, further confirming its relatively weak performance compared to other foundational and ML based methods.
Figure 6: Average ranks of forecasting methods with 95% confidence intervals based on the Nemenyi test for CRPS values. Lower ranks indicate better forecast performance.
5.2 Point and probabilistic forecast performances across forecast horizons
We also analyze the forecast performances over different horizons to evaluate how the methods perform over time. The forecast horizons range from month 1 to month 3, corresponding to the upcoming planning period used by planners for decision-making. First, we examine the error distribution across all methods. The RF method consistently shows the highest point forecast accuracy across all three horizons. Additionally, the top five methods, including the proposed hybrid weighted averaging method, maintain consistent performance throughout the forecast periods.
In terms of probabilistic forecast accuracy, similar patterns are observed across different methods. While these plots offer a high-level overview of error metric distributions (see Figure 10 in Appendix 1), they do not provide detailed insights into the differences between the top- and low-ranking methods. To gain a clearer understanding of the error metrics distribution, we plot density distributions, focusing on the top three and bottom three forecasting methods for both point forecasts and probabilistic forecasts.
Figure 7 and Figure 8 demonstrate that both the point and probabilistic forecast accuracy densities for the top three methods exhibit a narrower spread compared to the bottom three methods. This indicates that the forecast errors for these top methods are less variable and more consistently close to the actual values across different time series than those of the bottom three methods. The densities of all other methods, shown in grey, fall between those of the top and bottom methods, offering broader comparative context. Moreover, the plots show that the top methods maintain consistent performance across forecast horizons. However, it is noteworthy that the right tail of the density curves for RF and LGBM becomes more volatile as forecast errors increase, particularly at forecast horizon 3. This volatility may suggest that, while these top two methods often deliver consistently strong performance, there remain some uncertainties with specific time series that these methods are unable to capture effectively. In contrast, the Hybrid Weighted Averaging method shows a smoother tail, reflecting that it captures this variability more effectively compared to RF and LGBM.
Figure 7: The distribution of MASE values for the top three and bottom three forecasting methods across the horizons is presented. The methods are ranked based on their mean MASE values, with the top and bottom methods selected accordingly. Grey lines represent the distribution of MASE values for all other methods, providing a comparative context.
Figure 8: The distribution of CRPS values for the top three and bottom three forecasting methods across the horizons is presented. The methods are ranked based on their mean CRPS values, with the top and bottom methods selected accordingly. Grey lines represent the distribution of CRPS values for all other methods, providing a comparative context.
5.3 Forecast performance and computational efficiency
We now focus on the computational efficiency of the forecasting methods. In this study, computational efficiency is defined as the total runtime required for one iteration on the first rolling origin. The runtime was calculated based on this definition, and each method was retrained during each iteration. For this analysis, we focused solely on methods that generate both point and probabilistic forecasts from our candidate methods.
We used two environments: an R Studio local implementation on a device with an 11th Gen Intel(R) Core(TM) i5-1135G7 @ 2.40 GHz and 8 GB RAM, as well as Google Colab on both CPU and T4 GPU devices. To compute the runtime for combined statistical and ML methods, we averaged the runtime of the respective underlying methods. For the proposed hybrid methods, we added the runtime of the underlying methods to the time taken by the proposed method to combine the forecasts.
Table 5 shows that, although the RF is the best ranked method in Nemenyi test, it requires significantly more runtime compared to the other forecasting methods. On the other hand, TimeGPT stands out as the fastest method, outperforming all other methods in terms of runtime while still providing reasonable forecast accuracy. This is a notable exception, as it balances performance and computational efficiency well.
However, it is important to note that ML methods were trained using a normal CPU and one core due to technical challenges in the setup. With access to GPU devices or CPUs with multiple cores, we could likely improve the computational performance of these ML methods.
Figure 9 shows a clear relationship between runtime and accuracy improvement. Most of the top-performing methods fall into the moderate runtime category including the Hybrid Weighted Average method, with RF (the top ranked) being the slowest method. Interestingly, the Hybrid Bias Adjustment method also falls into the moderate runtime category but shows a relatively high accuracy error. However, the runtime of hybrid methods largely depends on the underlying methods selected for combination.
From a practical perspective, choosing the right method should balance both performance and runtime. It is a tradeoff between the extra computational cost incurred by more sophisticated methods that can handle uncertainties and the lower cost and simplicity of standard time series methods.
Table 5: Forecast performance and computational efficiency for each forecast method are ordered based on the mean MASE.
Method Mean MASE Mean CRPS Runtime (Minutes) Runtime type
RF 0.743 9.997 312.50 CPU
Hybrid weighted averaging 0.775 9.868 205.90 CPU with 4 cores
LGBM 0.833 10.131 153.50 CPU
ML combined 0.847 10.286 202.83 CPU
XGB 0.859 10.56 142.50 CPU
TimeGPT reg 1.258 15.635 0.45 Colab T4 GPU
MLR 1.269 12.611 10.27 CPU
TimeGPT 1.292 15.831 0.23 Colab T4 GPU
Chronos 1.305 15.018 30.08 Colab T4 GPU
BSTS reg 1.327 15.342 47.23 CPU with 4 cores
SBA 1.331 - 18.23 CPU with 4 cores
Moving average 1.373 15.701 5.19 CPU with 4 cores
Statistical combined 1.378 16.045 19.99 CPU with 4 cores
ETS 1.379 15.397 29.89 CPU with 4 cores
ARIMA 1.386 15.703 27.32 CPU with 4 cores
Lag Llama 1.483 15.84 39.39 Colab T4 GPU
BSTS demo 1.521 17.064 55.73 CPU with 4 cores
sNaive 1.603 17.511 17.56 CPU with 4 cores
Hybrid bias adjustment 4.360 29.062 209.40 CPU with 4 cores
Figure 9: Runtime vs. forecast performance (The X-axis shows the runtime speed for each method as fast, moderate, or slow).
6 Discussion
6.1 Findings
Among the methods evaluated, the Hybrid Weighted Averaging method stood out as a robust performer. In terms of mean MASE, it ranked second, and it achieved the top rank for mean CRPS, placing it on par with the best performing methods. Notably, the Nemenyi test revealed no significant performance differences between the Hybrid Weighted Averaging method and the RF method across both point and probabilistic forecasting. This result demonstrates that the Hybrid Weighted Averaging method is a reliable and accurate choice for forecasting contraceptive demand in contexts where probabilistic accuracy and low variance in forecasts are critical. Moreover, our analysis of forecast performance across multiple time horizons found that the Hybrid Weighted Averaging method maintained stable accuracy, which is crucial for demand planning.
One important limitation of the Hybrid Weighted Averaging method, however, is that it becomes less suitable when the point forecast deviates significantly from the central tendency of the probabilistic forecast. In such cases, the Hybrid Bias Adjustment method, designed to handle larger deviations, may be preferable. However, the bias adjustment method produced higher errors overall. In practice, this method can apply significant adjustments to the probabilistic forecast; therefore, obtaining expert opinion on its estimates would be beneficial for evaluating its performance more effectively.
The performance evaluation of both point and probabilistic forecasts across methods showed consistent results. The MASE and CRPS analyses reveal that top ML methods—RF, LGBM, XGB, and ML combined—consistently outperform time series methods in both point and probabilistic forecasting, with the RF method ranking highest on both metrics among ML methods. This suggests that ML methods generalize well across diverse time series patterns (smooth, erratic, lumpy, intermittent) within the FPSC context, effectively handling the underlying complexity of the data. Existing literature supports these findings, indicating that ML methods are better equipped to handle underlying uncertainties compared to time series methods [44]. Moreover, the consistent performance of ML methods underscores their robustness in capturing data dynamics over time. However, the MLR method did not perform as well as other ML methods. This discrepancy may stem from the linearity assumption inherent in MLR, whereas real-world FPSC data likely exhibit more complex, non-linear patterns, which MLR struggles to capture effectively.
Despite these results, time series methods should not be entirely discounted. For example, the SBA method, while outperformed by ML methods, surpassed many other time series approaches in terms of MASE, suggesting that it may be more suitable for site-level contraceptive demand data, which often exhibit low or zero demand. Syntetos et al. [68] also highlight the suitability of the SBA method for such scenarios, though its limitation lies in its inability to provide forecast distributions [34]. Additionally, it is notable that when BSTS methods were provided with time series based predictors and categorical and date features as regressors, they performed significantly better than when demographic predictors were used. A potential reason for this could be the annual granularity of demographic predictors, whereas this study focuses on monthly data.
Interestingly, foundational methods did not outperform top ML-based methods. When trained as univariate methods in a zero-shot setting, they performed similarly to time series methods, offering no clear advantage. Although foundational methods are typically trained on large time series datasets from various domains [25, 59, 5], the time series data observed in the FPSC context pose additional challenges such as noise, inaccuracy, and incompleteness [10]. This highlights the need for pretraining these methods on time series data from the humanitarian sector, which shares similar data challenges. Carriero et al. [15] found that foundational methods perform better with stationary time series, and they emphasize the importance of incorporating external factors such as expert knowledge. Our study corroborates this by showing that the incorporation of external regressors significantly improved the forecasting performance of the TimeGPT method.
Furthermore, the current methods applied to contraceptive demand planning demonstrate that both the moving average and demographic methods underperform, with the demographic method being the worst. One possible reason for this is that family planning indicators are often based on assumptions [2], and these indicators are calculated at the national or global level [50], making them less reflective of local patterns. Additionally, the demographic method provides estimates of total need rather than consumption, which can lead to discrepancies between estimates and actual consumption [2]. Additionally, our analysis of forecast performance across multiple time horizons revealed that method performance remained stable over time. Notably, the USAID competition submissions were not able to outperform the top five methods in our study.
Finally, we assessed the trade-off between computational efficiency and forecast accuracy. While RF achieved the high accuracy, it also demanded greater computational resources. The Hybrid Weighted Averaging, LGBM, and XGB methods offered a balanced solution, delivering high accuracy with moderate computational demands. time series, Bayesian, and foundational methods were computationally efficient but less accurate. TimeGPT with external regressors, though not the most accurate, achieved a balance between accuracy and efficiency, making it suitable for resource-constrained contexts where moderate accuracy is acceptable.
In practice, healthcare sites generate forecasts monthly, and thus, moderately efficient methods like LGBM are often a suitable choice. LGBM’s track record in forecasting competitions like M5 [44], as well as its strong performance in this study, further support its practical applicability. Accordingly, our proposed hybrid combination approach could be employed to combine judgment forecasts with probabilistic forecasts generated using LGBM.
6.2 Managerial implications
Demand forecasting for contraceptives in developing countries is a critical managerial task, given the volatile and unpredictable nature of demand. However, many field-level staff still rely on basic methods like moving averages or demographic projections, which often fall short in addressing these complexities. Our research underscores the need to transition to advanced probabilistic forecasting approaches that provide a range of potential outcomes rather than a single-point estimate. This shift can enable field-level staff to better anticipate demand variability and uncertainty. Thus, our proposed model can support procurement strategies by providing probabilistic forecasts that help optimize order quantities based on demand uncertainty. This enables field level staff to reduce both overstocking and understocking risks by quantifying the uncertainty. Additionally, the model enhances FPSC resilience by allowing inventory levels to be adjusted dynamically based on forecast distributions, improving responsiveness to demand fluctuations while ensuring consistent contraceptive availability.
Additionally, our findings highlight the importance of integrating domain expertise with ML forecasts to address the limitations of purely data-driven approaches. The variability in ML performance, particularly in capturing extreme demand patterns, points to the value of a hybrid approach. By allowing expert judgement to refine ML outputs, this method improves transparency and ensures alignment with the specific goals of FPSC. Field-level staff can therefore benefit from actionable insights while avoiding the “black box” nature of many advanced forecasting methods.
To further aid field-level staff, we developed a practical guideline (see Table LABEL:tbl-guide in Appendix 2) that compares various forecasting methods tested in our study. This resource builds on prior frameworks, such as the Contraceptive Forecasting Handbook by USAID [73], but goes further by incorporating advanced techniques like Bayesian modeling and hybrid methods tailored for uncertain demand environments. This guideline serves as a roadmap for field-level staff to select the most suitable method for their operational context, improving decision-making quality and efficiency.
Lastly, one of the broader implications of our study is its potential for replication in other sectors. The adaptability of our proposed method means that it can be applied to other humanitarian or public health contexts that deal with volatile demand patterns, such as food aid distribution or medical supply chains. The ability to generalize this approach across various sectors ensures that field-level staff in different industries can also benefit from improved forecasting practices, thereby increasing the overall reliability and resilience of their supply chains.
6.3 Limitations and future directions
While our study provides valuable insights into contraceptive demand forecasting using a variety of methods, certain limitations need to be acknowledged. First, although we compared the point and probabilistic forecast performance across different methods ranging from statistical, Bayesian, ML, to foundational, we did not conduct a detailed diagnostic analysis of how each method behaves in the presence of volatile time series, such as those typical of contraceptive demand. Volatile time series can exhibit erratic patterns, discontinuations, and unexpected spikes, complicating the forecasting process.
There are two main challenges in addressing this issue. First, methods trained in a global setting (where one method handles all series) allow for easier diagnostics. However, methods trained in a local setting (where one method is fitted per time series) make diagnostic processes significantly more complicated due to the large number of methods involved. Second, there is no standard diagnostic framework that applies across different model families, making it difficult to compare models with varied structures. Future research should explore the development of a standardized diagnostic framework for diverse forecasting models, particularly in the context of contraceptive demand, as such a framework could improve our understanding of how models behave under real-world complexities.
Another limitation is that our linear equal-weighted forecasts did not perform as well as expected. This may be due to the assumption that all forecasts were well-calibrated, and thus their combination would be too. However, the combined forecasts may have been miscalibrated, resulting in lower performance. This issue applies to our proposed methods as well. While some research on forecast calibration exists, such as the work by Ranjan and Gneiting [58], further investigation is needed to improve post-calibration processes in our hybrid methods and linear pooling approaches. Improving calibration could enhance both accuracy and reliability in demand forecasting.
On the other hand, FPSC is often subject to uncertainties arising from complex demand patterns, variable lead times, and dependence on donor support [49]. For instance, demographic factors like the age structure of a region can influence contraceptive demand. Haakenstad et al. [31] highlighted that young women (ages 15-25) tend to prefer short-term contraceptive methods, while older, married women are more likely to use long-term methods. However, even these behaviours are heavily influenced by social and cultural beliefs [62]. In some regions, such as India, long-term contraceptive methods are popular among younger women [32]. Infrastructure variability also poses challenges, particularly in rural and underserved areas, where logistical constraints can affect the distribution and accessibility of contraceptive products. Furthermore, Karimi et al. [38] highlight that rural facilities often face difficulties such as poor road conditions, inadequate storage, and delivery delays, further complicating supply chain operations. While field-level healthcare staff rely on their contextual knowledge to adjust forecasts, much of this information remains undocumented, making it difficult to systematically incorporate into forecasting models. Identifying and defining such influencing factors and integrating them into forecasting methods remains an important research area for future exploration.
Another key issue in the FPSC is the presence of censored demand due to stockouts, under-reporting, or discontinuations. In our modelling process, we did not account for these scenarios. Future research should explore how to develop forecasting methods that can handle stockout data, mitigate its impact on decision-making. Moreover, addressing the challenges of cold starts (multiple origin points) and cold ends (discontinuations) in time series forecasting is crucial, as these are prevalent in FPSC and should be considered in future methods.
Additionally, we did not consider product switching or substitution in response to availability or accessibility issues. Unlike other supply chains, contraceptive product substitution is challenging because each product has unique attributes, such as effectiveness and coverage period. Moreover, women’s preferences are influenced by health concerns—many women are reluctant to switch products they have used long-term due to perceived health risks [63]. Younger women, for instance, may avoid long-term contraceptives, fearing they could affect future fertility [32]. Investigating how to incorporate product-switching behaviors into the forecasting process is an important area for future research.
Another limitation of our study is we did not conduct an exploratory analysis of the sources of expert bias in judgemental forecasting. However, understanding when and why human forecasters introduce biases is crucial for improving hybrid forecasting accuracy. Future research could address this gap by systematically analysing the conditions under which expert bias occurs, particularly in the FPSC context. This could involve access to detailed expert forecasts, identifying systematic biases, and developing mechanisms to mitigate their impact. Such insights would further refine hybrid intelligence models by improving the integration of human intuition and algorithmic precision.
Finally, forecast distributions are just one aspect of logistics management in contraceptive demand forecasting. Decision-makers need to understand how to use forecast data for FPSC operations like inventory optimization, distribution, and procurement. As Raftery [57] suggests, forecasts may only need to provide prediction intervals or quantiles in some cases to inform decisions. Whether this approach applies to FPSC remains an open question. Future research should explore how to effectively communicate probabilistic forecasts and integrate them with inventory management, assessing the practical benefits for FPSC decision-making and improving planning and strategy formulation.
7 Conclusion
Effective forecasting and planning within the FPSC are essential to ensure that contraceptives are consistently and readily available to those who need them [49]. Accurate and reliable demand forecasting is therefore critical within the FPSC, as it supports informed decision-making to ensure access to safe and effective contraceptives. This, in turn, empowers individuals and communities to make informed reproductive health choices and helps reduce the unmet need for contraceptives [1].
Our study points out the need to improve contraceptive demand forecasting by combining probabilistic forecasting methods with expert knowledge, especially within the FPSC. Current forecasting methods often use simple methods, like moving averages or basic demographic approaches, which don’t fully capture the complexities of contraceptive demand. These patterns are influenced by various factors, including stockouts, product switching, and socio-demographic variables. While system-generated forecasts are good at showing past trends, literature shows that expert input is vital for refining forecasts in real-world situations with incomplete data and changing demand [22]. Therefore, we propose a new framework that enhances contraceptive demand forecasting by merging probabilistic methods with expert insights. This combined approach offers a promising solution for dealing with the uncertainties and complexities of contraceptive demand in developing countries.
Our proposed hybrid method, which combines point forecasts with probabilistic distributions, offers a promising way to improve forecasts by incorporating expert knowledge. The hybrid weighted averaging method strikes a good balance between accuracy and efficiency, making it effective for adjusting probabilistic forecasts where the algorithm has already accounted for most uncertainties. Although the hybrid bias adjustment method showed higher error rates, it allows for important adjustments to probabilistic forecasts using point forecasts, especially in situations with stockouts and incomplete data, offering greater flexibility to integrate expert judgment.
Furthermore, we review various forecasting methods, including time series, Bayesian, foundational time series, and machine learning methods, along with our new hybrid methods. We provide insights into the strengths and weaknesses of these methods, their computational efficiency, and their most appropriate use cases. This makes our study a useful guide for forecasting contraceptive demand.
In summary, our study addresses a key gap in probabilistic forecasting for contraceptive demand and presents a combined approach that blends algorithmic and human expertise. The findings from this study improve forecasting methods within the FPSC and offer practical recommendations for better contraceptive forecasting in developing countries.
Data availability statement
The R and Python code, along with the data required to reproduce all results in this paper, will be made available in a GitHub repository upon acceptance of the paper.
Acknowledgements
We would like to thank Laila Akhlaghi from John Snow, Inc. for her advice and feedback.
Disclosure statement
The authors report there are no competing interests to declare.
Appendix
Appendix 1
Figure 10: The first panel shows the distribution of MASE values for forecasting methods across different horizons. The boxplots are arranged in order of the median MASE values. The second panel shows the distribution of CRPS values for forecasting methods across different horizons. Similarly, the boxplots are arranged in order of the median CRPS values.
Appendix 2
Table 6: Guidelines for method selection in contraceptive demand forecasting.
Method Strengths Limitations Computational efficiency Suitable contexts Key assumptions
sNAÏVE Simple to implement, useful as a baseline forecast model Limited accuracy for non-stationary data, ignores trends and seasonality Very High (minimal computational cost) Benchmarking more advanced models, suitable for stable, short-term forecasting Assumes future demand will be exactly the same as the last observed period
Moving Average Smooths short-term fluctuations, useful for capturing general level trends Ignores seasonality, struggles with long-term trends, lags in response to sudden changes Very High (minimal computational cost) Suitable for stable demand with no major seasonality; underperforms in volatile environments Assumes future demand can be estimated by averaging past values within a chosen window
ETS Captures trend, seasonality, and noise; relatively easy to interpret Struggles with high volatility, assumes constant trend and seasonality High (fast computational time) Data with clear seasonality and trends, underperforms with volatile or intermittent demand data Assumes trend and seasonality are stable over time and can be modeled separately using smoothing
ARIMA Strong for univariate, stationary time series, handles seasonality well Requires stationarity, can struggle with high volatility, requires careful model tuning High (fast computational time) Suitable for stationary or seasonally adjusted data, poor in volatile or intermittent demand data Assumes data is stationary or can be made stationary through transformations (e.g., differencing)
SBA Designed specifically for intermittent demand Not suitable for continuous demand or high variability; does not handle probabilistic forecasting Very High (minimal computational cost) Effective for intermittent demand, with both frequent and infrequent zeroes, and with predictable inter-demand intervals Assumes demand occurs sporadically with zero demand periods, and uses probability-based predictions for inter-demand intervals
Multiple Linear Regression Easy to interpret, handles multiple predictors, including external factors Struggles with non-linearity, multicollinearity, and complex interactions Moderate (depends on number of predictors) Environments with clear, linear relationships between target variable and predictors Assumes linear relationships between the dependent and independent variables
LightGBM High accuracy in large, complex datasets, handles many types of predictors, efficient for large datasets Can overfit if not carefully tuned, sensitive to noise Moderate (more efficient than RF and XGBoost) High-dimensional data with complex, non-linear relationships Assumes non-linear and complex relationships that can be captured via gradient boosting algorithms
XGBoost High accuracy, robust to overfitting, handles complex interactions well Requires extensive tuning, computationally expensive compared to simpler models Moderate (more expensive than LightGBM) High-dimensional data, especially with complex relationships among predictors Assumes relationships between variables can be learned through gradient boosting with proper tuning
Random Forest High accuracy, handles non-linear patterns, robust for point/probabilistic forecasting Computationally expensive, can overfit on small datasets, slow for large datasets Low (slow, especially for large datasets) High-dimensional datasets with complex, non-linear relationships Assumes patterns in the data are driven by non-linear relationships learned via decision trees
Demographic Method Simple, interpretable, incorporates demographic factors, useful for forecasting in new product categories or long-term planning Poor handling of dynamic or volatile data, limited to demographic variables, not suitable for short-term forecasting Very High (minimal computational cost) Situations driven by demographic factors (e.g., population, age) or new product forecasting Assumes demographic factors like population size and age are primary drivers of demand
Bayesian Structural Time Series Captures seasonality, trends, and structural breaks; effective for small datasets Slow for complex models with many predictors; struggles in volatile environments Moderate (higher with more predictors) Data with clear seasonality, trends, or structural breaks Assumes seasonality, trends, and causal relationships can be captured through a Bayesian framework
TimeGPT (with Regressors) Highly computationally efficient, integrates external variables well Performance degrades without external regressors, limited for non-structured data Very High (minimal computational cost) Low-resource environments with strong external drivers (e.g., economic factors) Assumes external regressors are strongly correlated with demand patterns and that the time series follows stable patterns
Lag Llama Captures lag effects in demand, simple to implement Limited to contexts with strong lagged relationships; underperforms in complex scenarios High (higher computational cost) Situations with significant lag effects between past and future demand Assumes demand is heavily influenced by past values with strong lag effects, and future demand can be predicted by historical lags
Amazon Chronos Provides a strong baseline, simple to use Underperforms against advanced machine learning models; limited handling of external variables High (higher computational cost) Univariate time series forecasting with stationary or transformed data Assumes simple historical patterns can be extrapolated without requiring complex features or relationships
Hybrid Weighted Averaging Model Combines strengths of multiple models, stable across forecast horizons Sensitive to weight assignment, performance degrades with poor weight selection Moderate (depends on underlying models) Suitable for volatile or dynamic demand; can incorporate expert input for probabilistic forecasting Assumes that multiple models capture different aspects of the demand patterns and can be effectively weighted to improve forecasting accuracy
Hybrid Bias Adjustment Model Corrects systematic biases in statistical models, improves forecast accuracy Limited impact if biases are minimal, requires good bias detection Moderate (depends on underlying models) Ideal when systematic biases exist in forecast models; useful in dynamic or volatile demand environments Assumes that consistent, predictable biases exist in the base models and that they can be adjusted for better forecasting